the demand curve for a swiveled chair is given by p=4000(3-5) dollars per swiveled chair. where p is the price and q is the quantity , in thousand of swivel chair, demanded at that price. what quantity will.be demanded if the price per swivel chair is 256.60


1
Expert's answer
2021-11-08T19:37:43-0500

"256.40=4000*3^{-q}"

"3^{-q}=0.06415"

"-q\ln(3)=\ln(0.06415)"

"q=2.500."

At the price of $256.60 will be demanded 2500 chairs.